OpenStudy (anonymous):

WILL MEDAL Two rooms in a house are similar rectangles. Room A is 12 ft by 16 ft. The longer side of Room B is 4 ft shorter than twice the length of the shorter side of Room A. What are the dimensions of Room B? A. 15 × 20 B. 16 × 20 C. 16 × 16 D. 14 × 16

OpenStudy (anonymous):

x will be the longer side of room B and y will be the shorter of room B x = 2*12 - 4 = 24 - 4 =20 ft Since they are similar then they share a proportionality constant which is: 20/16 = 1.25 Now we should multiply the shorter side with 1.25: 12*1.25 = 15 So it is 15 by 20.